random sample of 12 size AA batteries for toys yield a mean of 3.85 hours with standard deviation, 1.15 hours.
(a) Find the critical value, t*, for a 99% CI. t* =
(b) Find the margin of error for a 99% CI.

Answer by Boreal(15235)

t (df=11,0.995)=3.106
half-interval is margin of error
t*s/sqrt(n)=3.106*1.15/sqrt(12)=1.03